5-2: Future Values 5-3: Present Values Problem Walk-Through Present and future values of a cash flow stream An investment will pay $50 at the end of each of the next 3 years, $250 at the end of Year 4, $350 at the end of Year 5, and $500 at the end of Year 6. a. If other investments of equal risk earn 11% annually, what is its present value? Round your answer to the nearest cent. b. If other investments of equal risk earn 11% annually, what is its future value? Round your answer to the nearest cent.You want to buy a car, and a local bank will lend you $15,000. The loan will be fully amortized over 5 years (60 months), and the nominal interest rate will be 8% with interest paid monthly. a. What will be the monthly loan payment? Round your answer to the nearest cent. b. What will be the loans EAR? Round your answer to two decimal places.
Answer #1

We can solve this using excel NPV function:

Year Cash Flow
1 $    50.00
2 $    50.00
3 $    50.00
4 $ 250.00
5 $ 350.00
6 $ 500.00
PV $761.90

NPV formula:
=NPV(0.11,I9:I14)

Book1 (version 1) - AutoRecovered - Excel View Add-ins HelpTell me what you want to do Currency Imran Khan File Home Insert Page Layout Formulas Revie share comments Σ AutoSum . A De Copy Merge & Center . | $ . % , Conditional Format as Cell Formatting Table Styles nsert Delete Format Format Painter Sort & Find & Filter-Select Clear- Clipboard Alignment Number Cells Editing x fNPV(0.11,19:114) Year Cash Flow 1$ 50.00 2 $ 50.00 3$ 50.00 $250.00 5$350.00 6 $500.00 16 114 17 Sheet1Sheet2 Sheet3 3:48 AM O Type here to search 1/16/2019 21)

Future Value = Present Value*((1+r)^n)

r is the rate and n is the number of periods

Future value= 761.90*(1.11^6)=$1,425.06

PV 15000
n 60
I/Y 0.67%
FV 0
CPT PMT $304.15

EAR=((1.0067^12)-1)%
=8.30%